Helium in the amount of 3 moles expands in the process 1-2. The temperature of helium in state 1 is 273 K and in state 2 it is 373 K. How much heat is delivered to helium in the process 1-2? Universal gas constant is 8, 31 J/(mol • K). Give the answer in kJ and round it up to tenths.
1
Expert's answer
2020-11-12T09:37:10-0500
"Q=\frac{5}{2}nR(T_f-T_i)\\Q=\frac{5}{2}(3)(8.31)(373-273)=6232.5\ J"